OpenStudy (anonymous):

At 2 p.m. the shadow of a lighthouse is 19 feet long and the angle of elevation is 75*. Find the height of the lighthouse. Round to the nearest tenth. A. 70.9 ft B. 72.6 ft C. 21.1 ft D. 154.7 ft

OpenStudy (anonymous):

 tan(75*)=h/19 h=19tan75* h=70.9 so A.